OpenStudy (anonymous):

what's this?( in algebraic properties): (2 + x) + y = 2 + (x + y)

OpenStudy (anonymous):

well what properties do you know? :) try it first

OpenStudy (anonymous):

is it ditributive? @robtobey

OpenStudy (anonymous):

no distributive is a(b+c) = a*b + a*c for example. what else. also. distributive is only for multiplication. addition has its own properties

OpenStudy (anonymous):

so is it Transitive Property ?

OpenStudy (mertsj):

I see that on the left the x is paired with the 2. On the right, the x is paired with the y. What property deals with pairing or groupint?

OpenStudy (mertsj):

grouping.

OpenStudy (mertsj):

(a+b)+c=a+ (b+c) Look at your properties and see which one it is.

OpenStudy (anonymous):

communitve? @Mertsj

OpenStudy (anonymous):

@Mertsj ... is it communitive?

OpenStudy (mertsj):

noi. The commutative property says that you can change the order: a+b = b+a